What is the value of x? You are given x + 4 and 3x + 4.
Answer by mmm4444bot(95)

Hello There:
Before you can find the value of x, you need to have an equation.
An equation always contains an equals sign. What you have instead are two "expressions".
If you meant to type: x + 4 = 3x + 4
then you start by moving all of the terms that contain the variable x to one side of the equals sign, and by moving all of the constants to the other side.
We can move the x on the left side by subtracting x from both sides of the equation:
x - x + 4 = 3x - x + 4
4 = 2x + 4
Now we want to isolate the 2x, so we move the constant 4 from the right side to the left side. We do this by subtracting 4 from both sides.
4 - 4 = 2x + 4 - 4
0 = 2x
This last equation clearly shows that x must be zero.
If this is not the answer you are looking for, then please repost your question and be sure to provide all of the information so that we know what is equal to what.
~ Mark